OpenStudy (anonymous):

Solve 2 cos x + sec x = 3 on the interval [0,2pi)

OpenStudy (p0sitr0n):

recall secx=1/cosx 2cosx +1/cosx =3 2cos^2x +1 = 3cosx 2cos^2x-3cosx +1=0 Now the trick is to take y=cosx 2y^2-3y+1=0 solve this quadratic, then come back to y=cosx and reverse the roots you found , but keep the answers on [0,2pi]